one gallon of gasoline in an automobile's engine produces on the average 9.5 kg of carbon dioxide, which is a greenhouse gas, that is, it promotes the warming of Earth's atmosphore. Calculate the annual production of carbon dioxide in kilograms if there are 250 million cars in the United States and each car covers a distance of 5000 mi at a consumption rate of 20 miles per gallon.


1
Expert's answer
2021-10-20T02:52:05-0400

A gallon gasoline → 9.5 kg CO2

250 000 000 cars

1 car → 5000 miles

"250 \times 10^6 \times 5000 = 12500 \times 10^9 \;miles"

Total distance by all cars "= 12500 \times 10^9 \;miles"

20 miles → 1 gallon

"12500 \times 10^9 \;miles \; require = \frac{12500 \times 10^9}{20} = 62.5 \times 10^9 \;gallons"

Amount of gasoline used "= 62.5 \times 10^9 \;gallons"

Total amount of CO2 production

"= 62.5 \times 10^9 \times 9.5 = 593.75 \times 10^9 \;kg \;CO_2"

Answer: "593.75 \times 10^9 \;kg"